Circular No. 9070

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 Q5 (McNAUGHT) R. H. McNaught reports his discovery of a comet with a circular 18′′

coma on CCD images obtained with the 0.5-m Uppsala Schmidt telescope (discovery tabulated below); five stacked 30-s images obtained after moonset on Aug. 31.7–31.8 UT show a 0′.4 coma and a 50′′ tail in p.a. 285^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have noted the object’s cometary appearance. S. Foglia reports that fifteen stacked 60-s CCD images taken by R. Holmes (Charleston, IL, U.S.A., 0.61-m telescope) reveal a round coma of diameter 10′′ with a 30′′ tail in p.a. 260^{o}. Fifteen stacked 120-s unfiltered exposures taken remotely on Sept. 1.3 by E. Guido and G. Sostero with a 0.25-m reflector near Mayhill, NM, U.S.A., show a compact coma of diameter ∼ 12′′ and a broad tail nearly 20′′ long in p.a. 255^{o}. H. Sato (Tokyo, Japan, remotely using a Rent-A-Scope 0.25-m reflector near Mayhill on Sept. 1.35) finds a diffuse coma of diameter 8′′–20′′ and no tail. Frames taken by R. Miles with the Haleakala 2.0-m ‘Faulkes Telescope North’ on Sept. 1.6 show a diffuse 15′′ coma elongated toward p.a. 285^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Aug. 31.63470 1^{h}29^{m}04.^{s}36 −17^{o}26′52′′.9 17.0 McNaught

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-R02.

T = 2010 Mar. 11.218 TT  ω = 293.^{o}151
Ω = 160.070 }2000.0
q = 1.55732 AU  i = 40.823

COMETS C/2009 D7 AND C/2009 D8 (SOHO) Further to IAUC 9059, additional presumed comets have been found

on SOHO website images. C/2009 D7 was bright (mag ∼ 4) with a short tail in LASCO C3 images, while it showed a short tail and faded quickly in C2 images; this comet was also detected in STEREO COR2-A and COR2-B images with a tail ∼ 0^{o}.3 long. C/2009 D8 was slightly elongated, small, and stellar in apperance (mag ∼ 7, but it faded rapidly).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 D7 Feb. 26.029 23^{h}09.^{m}7 −10^{o}31′ C3/2* MU 2009-N01
C/2009 D8  26.129 22 42.2 − 7 03 C2  BZ 2009-N01
2009 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green